OpenStudy (anonymous):

. Solve: 6x - 10x + 5 = -15 A. x = 5 B. x = -15 C. x = 2.5 D. x = -5

OpenStudy (anonymous):

6x - 10x + 5 = -15 -4x + 5 = -15 -4x = -20 x = 5

OpenStudy (anonymous):

hold on im figuring it out

OpenStudy (anonymous):

hehe, too late, next time do it then post it :P

OpenStudy (anonymous):

its fine i got the same thing

OpenStudy (anonymous):

i had to go over it once but then came up with all the same

OpenStudy (anonymous):

you also can put the question and put what you did, and i will be more than happy to show you what you did wrong or if it is right :D

OpenStudy (anonymous):

thank you for the help